Just re-checked across C1, 0ohms.
Added a 104 cap where C5 goes, and the BPF moved to an "8" and stayed there. Did not change at all in reaction to trimmer setting with the extra cap in place.
Remelted the solder on the volume pot and double checked connections, so it is fully connected now if it was not before(I didn't check this one prior). Afterwards was able to at least to set the 3 nulls to their minimums.
I will need to wait till the band gets more traffic in the morning to see if I can receive anything.